# A Test of Substellar Evolutionary Models with High-Precision Ages from Asteroseismology and Gyrochronology for the Benchmark System HR 7672AB

### v1

@@ 3.2 Rotational Modeling
In addition to this physics-based approach, which models stellar rotation using an angular momentum loss prescription, we also consider empirical rotation-age-temperature relations, which are more data-driven. We adopted two empirical models. The gpgyro model (Lu2024) calibrates the spin-down relation using stellar clusters and kinematic ages. For this star, it predicts an age of $2.58\pm 0.17$ Gyr, using $P_{\rm rot}$, $T_{\text{eff}}$, and the absolute Gaia G-band magnitude as inputs. The gyrointerp model (Bouma2024) is calibrated on stellar clusters and predicts an age of $2.64\pm 0.51$ Gyr, based on $P_{\rm rot}$ and $T_{\text{eff}}$. The larger uncertainty in the latter estimate likely reflects a more realistic treatment of the intrinsic spread in gyrochrones, with considerations for the empirical limits of gyrochronology.
These rotation-based age estimates are listed in Table 7. A comparison of all estimates is shown in Figure 4, where we find that the results are broadly consistent across methods. However, the rotation-based ages tend to be systematically higher than those derived from asteroseismology, although the difference is not statistically significant. This offset could arise from systematic effects in either method. For example, latitudinal differential rotation can limit the accuracy of rotation-based estimates (Epstein2014), while uncertainties in the initial helium abundance can affect the seismic diagnostics of the sound-speed profile. To mitigate such systematics, a more robust approach is to perform a joint modeling of both methods, which we explore next.
3.3 Asteroseismic and Rotational Joint Modeling
@@
In this construction, we treat each class of observational constraint independently, so we do not expect any overconstraining of the age due to double-counting of information. We obtained a final age estimate of $t_{\star}=2.26\pm 0.40$ Gyr. The joint estimate falls between the asteroseismic and rotational ages, with an uncertainty of 18%. These results are summarized in Table 7 and shown in Figure 4. For comparison, Crepp2012 estimated the system age to be $2.5\pm 1.8$ Gyr from isochrone fitting and $2.5\pm 0.7$ Gyr from gyrochronology relations using mam08-ages. The combination of asteroseismology and gyrochronology employed in this work relies on independent diagnostics, making the resulting age scale more accurate and less affected by method-specific systematics. We also find an improvement in precision compared to previous estimates.
4 Testing Brown Dwarf Cooling Models
@@
Our new high-precision measurements of the HR 7672B dynamical mass and the age of the HR 7627AB system, combined with the existing HR 7672B luminosity measurement, enable this benchmark system to test substellar evolutionary models with a degree of rigor previously not possible. We consider six widely-used model grids: Burrows1997, Saumon2008, Baraffe2015, Phillips2020, Chabrier2023, and Morley2024. The different models employ different interior physics (e.g., equations of state) and boundary conditions (e.g., atmosphere models). Brown-dwarf evolution grids have progressed from early non-gray, baseline cooling tracks to modern models that fold in updated H/He equations of state (EOS), opacities, cloud physics, chemistry, and vertical mixing. Burrows1997 (Burrows1997) sets the foundation with non-gray atmospheres and coupled spectra/cooling. SM2008 (Saumon2008) introduced a hybrid cloudy/clear prescription to capture the L/T transition. BHAC2015 (Baraffe2015) refreshed pre-main-sequence/young-BD tracks with boundary conditions based on BT-Settl model atmospheres. ATMO2020 (Phillips2020) added a new H-He EOS and expanded opacities and chemistry. CBPD2023 (Chabrier2023) then isolated interior updates with a newer EOS, shifting hydrogen-burning limits and late-time cooling. Diamondback delivered a modern cloudy grid for warm L/T objects that systematically varies metallicity and cloud sedimentation. We use the solar-metallicity versions of all these models, in accord with the metallicity of HR 7672A.
@@
Along our new dynamical mass ($M=75.39\pm 0.67$ $\text{M}_{\text{Jup}}$) and host star age ($t=2.26\pm 0.40$ Gyr), the third observation of HR 7672B needed to test models is the bolometric luminosity. We adopt $\log(L/{\text{L}_{\odot}})=-4.19\pm 0.04$ dex from Brandt2019, which represents the weighted average of values derived from the companion’s $H$-band ($4.23\pm 0.05$ dex) and $K_{S}$-band ($4.14\pm 0.06$ dex) photometry measured by Boccaletti2003 and then combined with a bolometric correction from Dupuy2017. We note that the Boccaletti2003 photometry leads to a color of $(H-K_{S})=1.00\pm 0.17$ mag, which is somewhat redder compare to colors for mid-L dwarfs, e.g., $(H-K_{S})=0.64_{-0.20}^{+0.11}$ mag from the L4–L4.9 template of 2018ApJS..234....1B.
@@
$\beta$ Pic c 86.8212, -51.0665  $\cdots$ A6V 0.13″ (2.6 au) $8.3\pm 1$ $4.43\pm 0.10$jjUsing the 2023ApJ...959...63S relation between $M(K_{MKO})$ and $\log(\mbox{$L_{bol}$})$ for young objects, we compute the difference in $\log(\mbox{$L_{bol}$})$ between $\beta$ Pic b and $\beta$ Pic c, and then use the value for the former from Kamm25 to compute the latter. Photometry errors are propagated in the calculation, as is the intrinsic 0.057 dex scatter in the Sanghi et al. relation. $0.024\pm 0.003$ Lagr19;Nowa20; –; Bran21c; Nowa20, Li26; Bell15
@@
HD 63754 Bdd(1) Object: SIMBAD indicates this companion is improperly designated as HD 63754 B in 2024MNRAS.533.3501L, since the Washington Double Star catalog (2001AJ....122.3466M) notes a faint ($\Delta{\rm mag}\approx 7$) companion at $\approx$5″ separation (known as **SEE 87B), first observed by Thomas J. J. See in 1897. 2014AJ....147...86T note that the companion does not share the same proper motion as the primary (a.k.a. HIP 38216) based on an observation in 2016. However, Gaia DR3 (GaiaCollaboration2023) show the companion shares a common parallax with the primary, and thus the designation for the ultracool companion should be changed to ”C”. (2) Host Age: The age is reported by Li24 as >>3.4 Gyr at 68% CL and 2.4 Gyr at >>95% CL. We report this here as 2.4-10 Gyr, but this is not a particularly good description of the posterior (see Figure 1 of 2024MNRAS.533.3501L). 117.4386, -20.2076  $\cdots$ G0V 0.47″ (23.6 au) $81.9_{-5.8}^{+6.4}$ $4.55\pm 0.08$ 2.4-10dd(1) Object: SIMBAD indicates this companion is improperly designated as HD 63754 B in 2024MNRAS.533.3501L, since the Washington Double Star catalog (2001AJ....122.3466M) notes a faint ($\Delta{\rm mag}\approx 7$) companion at $\approx$5″ separation (known as **SEE 87B), first observed by Thomas J. J. See in 1897. 2014AJ....147...86T note that the companion does not share the same proper motion as the primary (a.k.a. HIP 38216) based on an observation in 2016. However, Gaia DR3 (GaiaCollaboration2023) show the companion shares a common parallax with the primary, and thus the designation for the ultracool companion should be changed to ”C”. (2) Host Age: The age is reported by Li24 as >>3.4 Gyr at 68% CL and 2.4 Gyr at >>95% CL. We report this here as 2.4-10 Gyr, but this is not a particularly good description of the posterior (see Figure 1 of 2024MNRAS.533.3501L). Li24; –; Li24; Li24; Li24
@@
Gl 417B 168.1068, +35.8035 L4.5: G2V 90″ (2,133.0 au) $58.2\pm 1.5$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $4.132_{-0.031}^{+0.029}$ $0.75_{-0.12}^{+0.15}$ Kirk00; Dupu12; Dupu17, Li26; Dupu17; Dupu14
Gl 417C 168.1068, +35.8035 L6: G2V 90″ (2,133.0 au) $51.8\pm 2.4$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $4.22\pm 0.03$ $0.75_{-0.12}^{+0.15}$ Kirk00; Dupu12; Dupu17, Li26; Dupu17; Dupu14
@@
HD 130948 B 222.5667, +23.9116 L4: F9IV-V 2.64″ (48.0 au) $59.4\pm 0.9$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $3.85_{-0.06}^{+0.07}$ $0.79_{-0.15}^{+0.22}$eeComputed from gyrochronology, where the posterior of the inferred age distribution is normal in $\log$(age). Pott02; Goto02; Dupu17, Li26; Dupu17; Dupu09a
HD 130948 C 222.5667, +23.9116 L4: F9IV-V 2.64″ (48.0 au) $56.5\pm 1.8$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $3.96\pm 0.06$ $0.79_{-0.15}^{+0.22}$eeComputed from gyrochronology, where the posterior of the inferred age distribution is normal in $\log$(age). Pott02; Goto02; Dupu17, Li26; Dupu17; Dupu09a
Gl 569 Ba 223.6226, +16.1024 M8.5 M3V 5″ (49.7 au) $80.1_{-8}^{+8.6}$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $3.44\pm 0.03  \cdots$ Skru87b; Henr90;Lane01; Kono10, Dupu17; Dupu17; –
Gl 569 Bb 223.6226, +16.1024 M9.0 M3V 5″ (49.7 au) $57.5_{-8.5}^{+6.8}$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from Dupuy2017 have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in 2009ApJ...692..729D and 2014ApJ...790..133D, respectively. $3.67\pm 0.05  \cdots$ Skru87b; Henr90;Lane01; Kono10, Dupu17; Dupu17; –
@@
Gl 802 B 310.8303, +55.3481  $\cdots$ M5Ve 0.09″ (1.5 au) $66\pm 5$ $4.43\pm 0.06$ggLuminosity computed from $M(K)$ in 2008ApJ...678..463I and the $M(K)-\mbox{$L_{bol}$}$ relation in 2017ApJS..231...15D. 8-11 Prav05; –; Irel08; Li26; Belo18, Helm18
@@
HD 206893 B 326.3413, -12.7834 L6+/-2 F5V 0.27″ (11.0 au) $28_{-2.1}^{+2.2}$ $4.24_{-0.09}^{+0.1}$hh(1) Log($L_{bol}$): 2023A&A...671L...5H report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. $0.25_{-0.2}^{+0.45}$hh(1) Log($L_{bol}$): 2023A&A...671L...5H report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. Mill17; Ward21; Hink23; Kamm25; Delo17b
HD 206893 c 326.3413, -12.7834  $\cdots$ F5V 0.11″ (4.5 au) $12.7_{-1}^{+1.2}$ $4.42_{-0.1}^{+0.12}$ $0.25_{-0.2}^{+0.45}$hh(1) Log($L_{bol}$): 2023A&A...671L...5H report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. Hink23; –; Hink23; Kamm25; Delo17b
@@
Note. — In addition, 2021ApJ...913L..26B discovered a faint companion to the old ($\approx$7–10 Gyr) G5V star HD 417127, with a dynamical mass of 67.5–177 $\text{M}_{\text{Jup}}$ (95% confidence limits) measured by combining RV, imaging, and Hipparcos+Gaia astrometry. The companion’s $H$-band absolute magnitude is too faint relative to this measured mass, suggesting the companion may be an unresolved binary.

### v2

@@ III.2 Rotational Modeling
In addition to this physics-based approach, which models stellar rotation using an angular momentum loss prescription, we also consider empirical rotation-age-temperature relations, which are more data-driven. We adopted two empirical models. The gpgyro model (Y. Lu et al., 2024) calibrates the spin-down relation using stellar clusters and kinematic ages. For this star, it predicts an age of $2.58\pm 0.17$ Gyr, using $P_{\rm rot}$, $T_{\text{eff}}$, and the absolute Gaia G-band magnitude as inputs. The gyrointerp model (L. G. Bouma et al., 2024) is calibrated on stellar clusters and predicts an age of $2.64\pm 0.51$ Gyr, based on $P_{\rm rot}$ and $T_{\text{eff}}$. The larger uncertainty in the latter estimate likely reflects a more realistic treatment of the intrinsic spread in gyrochrones, with considerations for the empirical limits of gyrochronology.
These rotation-based age estimates are listed in Table 7. A comparison of all estimates is shown in Figure 8, where we find that the results are broadly consistent across methods. However, the rotation-based ages tend to be systematically higher than those derived from asteroseismology, although the difference is not statistically significant. This offset could arise from systematic effects in either method. For example, latitudinal differential rotation can limit the accuracy of rotation-based estimates (C. R. Epstein & M. H. Pinsonneault, 2014), while uncertainties in the initial helium abundance can affect the seismic diagnostics of the sound-speed profile. To mitigate such systematics, a more robust approach is to perform a joint modeling of both methods, which we explore next.
III.3 Asteroseismic and Rotational Joint Modeling
@@
In this construction, we treat each class of observational constraint independently, so we do not expect any overconstraining of the age due to double-counting of information. We obtained a final age estimate of $t_{\star}=2.26\pm 0.40$ Gyr. The joint estimate falls between the asteroseismic and rotational ages, with an uncertainty of 18%. These results are summarized in Table 7 and shown in Figure 8. For comparison, J. R. Crepp et al. (2012) estimated the system age to be $2.5\pm 1.8$ Gyr from isochrone fitting and $2.5\pm 0.7$ Gyr from gyrochronology relations using E. E. Mamajek & L. A. Hillenbrand (2008). The combination of asteroseismology and gyrochronology employed in this work relies on independent diagnostics, making the resulting age scale more accurate and less affected by method-specific systematics. We also find an improvement in precision compared to previous estimates.
IV Testing Brown Dwarf Cooling Models
@@
Our new high-precision measurements of the HR 7672B dynamical mass and the age of the HR 7627AB system, combined with the existing HR 7672B luminosity measurement, enable this benchmark system to test substellar evolutionary models with a degree of rigor previously not possible. We consider six widely-used model grids: A. Burrows et al. (1997, Burrows1997), D. Saumon & M. S. Marley (2008, SM2008), I. Baraffe et al. (2015, BHAC2015), M. W. Phillips et al. (2020b, ATMO2020), G. Chabrier et al. (2023, CBPD2023), and C. V. Morley et al. (2024, Diamondback). The different models employ different interior physics (e.g., equations of state) and boundary conditions (e.g., atmosphere models). Brown-dwarf evolution grids have progressed from early non-gray, baseline cooling tracks to modern models that fold in updated H/He equations of state (EOS), opacities, cloud physics, chemistry, and vertical mixing. Burrows1997 (A. Burrows et al., 1997) sets the foundation with non-gray atmospheres and coupled spectra/cooling. SM2008 (D. Saumon & M. S. Marley, 2008) introduced a hybrid cloudy/clear prescription to capture the L/T transition. BHAC2015 (I. Baraffe et al., 2015) refreshed pre-main-sequence/young-BD tracks with boundary conditions based on BT-Settl model atmospheres. ATMO2020 (M. W. Phillips et al., 2020b) added a new H-He EOS and expanded opacities and chemistry. CBPD2023 (G. Chabrier et al., 2023) then isolated interior updates with a newer EOS, shifting hydrogen-burning limits and late-time cooling. Diamondback delivered a modern cloudy grid for warm L/T objects that systematically varies metallicity and cloud sedimentation. We use the solar-metallicity versions of all these models, in accord with the metallicity of HR 7672A.
@@
Along our new dynamical mass ($M=75.39\pm 0.67$ $\text{M}_{\text{Jup}}$) and host star age ($t=2.26\pm 0.40$ Gyr), the third observation of HR 7672B needed to test models is the bolometric luminosity. We adopt $\log(L/{\text{L}_{\odot}})=-4.19\pm 0.04$ dex from T. D. Brandt et al. (2019a), which represents the weighted average of values derived from the companion’s $H$-band ($4.23\pm 0.05$ dex) and $K_{S}$-band ($4.14\pm 0.06$ dex) photometry measured by A. Boccaletti et al. (2003) and then combined with a bolometric correction from T. J. Dupuy & M. C. Liu (2017a). We note that the A. Boccaletti et al. (2003) photometry leads to a color of $(H-K_{S})=1.00\pm 0.17$ mag, which is somewhat redder compare to colors for mid-L dwarfs, e.g., $(H-K_{S})=0.64_{-0.20}^{+0.11}$ mag from the L4–L4.9 template of W. M. J. Best et al. (2018).
@@
$\beta$ Pic c 86.8212, -51.0665  $\cdots$ A6V 0.13″ (2.6 au) $8.3\pm 1$ $4.43\pm 0.10$jjUsing the A. Sanghi et al. (2023) relation between $M(K_{MKO})$ and $\log(\mbox{$L_{bol}$})$ for young objects, we compute the difference in $\log(\mbox{$L_{bol}$})$ between $\beta$ Pic b and $\beta$ Pic c, and then use the value for the former from Kamm25 to compute the latter. Photometry errors are propagated in the calculation, as is the intrinsic 0.057 dex scatter in the Sanghi et al. relation. $0.024\pm 0.003$ Lagr19;Nowa20; –; Bran21c; Nowa20, Li26; Bell15
@@
HD 63754 Bdd(1) Object: SIMBAD indicates this companion is improperly designated as HD 63754 B in Y. Li et al. (2024), since the Washington Double Star catalog (B. D. Mason et al., 2001) notes a faint ($\Delta{\rm mag}\approx 7$) companion at $\approx$5″ separation (known as **SEE 87B), first observed by Thomas J. J. See in 1897. A. Tokovinin (2014) note that the companion does not share the same proper motion as the primary (a.k.a. HIP 38216) based on an observation in 2016. However, Gaia DR3 (Gaia Collaboration et al., 2023) show the companion shares a common parallax with the primary, and thus the designation for the ultracool companion should be changed to ”C”. (2) Host Age: The age is reported by Li24 as >>3.4 Gyr at 68% CL and 2.4 Gyr at >>95% CL. We report this here as 2.4-10 Gyr, but this is not a particularly good description of the posterior (see Figure 1 of Y. Li et al., 2024). 117.4386, -20.2076  $\cdots$ G0V 0.47″ (23.6 au) $81.9_{-5.8}^{+6.4}$ $4.55\pm 0.08$ 2.4-10dd(1) Object: SIMBAD indicates this companion is improperly designated as HD 63754 B in Y. Li et al. (2024), since the Washington Double Star catalog (B. D. Mason et al., 2001) notes a faint ($\Delta{\rm mag}\approx 7$) companion at $\approx$5″ separation (known as **SEE 87B), first observed by Thomas J. J. See in 1897. A. Tokovinin (2014) note that the companion does not share the same proper motion as the primary (a.k.a. HIP 38216) based on an observation in 2016. However, Gaia DR3 (Gaia Collaboration et al., 2023) show the companion shares a common parallax with the primary, and thus the designation for the ultracool companion should be changed to ”C”. (2) Host Age: The age is reported by Li24 as >>3.4 Gyr at 68% CL and 2.4 Gyr at >>95% CL. We report this here as 2.4-10 Gyr, but this is not a particularly good description of the posterior (see Figure 1 of Y. Li et al., 2024). Li24; –; Li24; Li24; Li24
@@
HIP 54515 B 167.3069, -1.3333 L0+/-2 A5V 0.14″ (11.6 au) $17.7_{-4.9}^{+7.6}$ $3.52\pm 0.03$ $0.115_{-0.092}^{+0.085}$ Curr25; Curr25; Curr25; Curr25; Curr25
Gl 417B 168.1068, +35.8035 L4.5: G2V 90″ (2,133.0 au) $58.2\pm 1.5$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $4.132_{-0.031}^{+0.029}$ $0.75_{-0.12}^{+0.15}$ Kirk00; Dupu12; Dupu17, Li26; Dupu17; Dupu14
Gl 417C 168.1068, +35.8035 L6: G2V 90″ (2,133.0 au) $51.8\pm 2.4$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $4.22\pm 0.03$ $0.75_{-0.12}^{+0.15}$ Kirk00; Dupu12; Dupu17, Li26; Dupu17; Dupu14
@@
HD 130948 B 222.5667, +23.9116 L4: F9IV-V 2.64″ (48.0 au) $59.4\pm 0.9$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $3.85_{-0.06}^{+0.07}$ $0.79_{-0.15}^{+0.22}$eeComputed from gyrochronology, where the posterior of the inferred age distribution is normal in $\log$(age). Pott02; Goto02; Dupu17, Li26; Dupu17; Dupu09a
HD 130948 C 222.5667, +23.9116 L4: F9IV-V 2.64″ (48.0 au) $56.5\pm 1.8$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $3.96\pm 0.06$ $0.79_{-0.15}^{+0.22}$eeComputed from gyrochronology, where the posterior of the inferred age distribution is normal in $\log$(age). Pott02; Goto02; Dupu17, Li26; Dupu17; Dupu09a
Gl 569 Ba 223.6226, +16.1024 M8.5 M3V 5″ (49.7 au) $80.1_{-8}^{+8.6}$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $3.44\pm 0.03  \cdots$ Skru87b; Henr90;Lane01; Kono10, Dupu17; Dupu17; –
Gl 569 Bb 223.6226, +16.1024 M9.0 M3V 5″ (49.7 au) $57.5_{-8.5}^{+6.8}$ccIndividual component masses are computed from the total dynamical mass of the binary, the flux ratio, and an evolutionary model. While strictly speaking these results are model-dependent, the near-equal flux ratios suggest that systematics from this process are small. For HD 130948BC and GJ 417BC, the results from T. J. Dupuy & M. C. Liu (2017a) have been updated to account for the change in parallax from Hipparcos to Gaia DR3, using the mass ratios estimated in T. J. Dupuy et al. (2009) and T. J. Dupuy et al. (2014), respectively. $3.67\pm 0.05  \cdots$ Skru87b; Henr90;Lane01; Kono10, Dupu17; Dupu17; –
@@
Gl 802 B 310.8303, +55.3481  $\cdots$ M5Ve 0.09″ (1.5 au) $66\pm 5$ $4.43\pm 0.06$ggLuminosity computed from $M(K)$ in M. J. Ireland et al. (2008) and the $M(K)-\mbox{$L_{bol}$}$ relation in T. J. Dupuy & M. C. Liu (2017b). 8-11 Prav05; –; Irel08; Li26; Belo18, Helm18
@@
HD 206893 B 326.3413, -12.7834 L6+/-2 F5V 0.27″ (11.0 au) $28_{-2.1}^{+2.2}$ $4.24_{-0.09}^{+0.1}$hh(1) Log($L_{bol}$): S. Hinkley et al. (2023) report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. $0.25_{-0.2}^{+0.45}$hh(1) Log($L_{bol}$): S. Hinkley et al. (2023) report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. Mill17; Ward21; Hink23; Kamm25; Delo17b
HD 206893 c 326.3413, -12.7834  $\cdots$ F5V 0.11″ (4.5 au) $12.7_{-1}^{+1.2}$ $4.42_{-0.1}^{+0.12}$ $0.25_{-0.2}^{+0.45}$hh(1) Log($L_{bol}$): S. Hinkley et al. (2023) report a more precise value of $4.23\pm 0.01$ dex, but this comes from fitting only a single set of model atmospheres, whereas Kamm25 fit multiple sets and also add in quadrature a 10% systematic uncertainty due to flux calibration errors. (2) Host Age: The reported error bars on the Delo17b ages ”should be considered as very conservative (max allowed range)”. Hink23; –; Hink23; Kamm25; Delo17b
@@
Note. — In addition, B. P. Bowler et al. (2021) discovered a faint companion to the old ($\approx$7–10 Gyr) G5V star HD 417127, with a dynamical mass of 67.5–177 $\text{M}_{\text{Jup}}$ (95% confidence limits) measured by combining RV, imaging, and Hipparcos+Gaia astrometry. The companion’s $H$-band absolute magnitude is too faint relative to this measured mass, suggesting the companion may be an unresolved binary.